# spacing=nonfrench

### 04 Mar 2026 v1

@@ 2.3. The universal $f^{\circ r}$-consistent ring
Denote by $r$ the residue degree of $L/\mathbb{Q}_{p}$, so that $q=p^{r}$. To suit this parameter, we adapt the universal $f$-consistent ring defined in [Spe18] by replacing $f$ with its $r$-th iterate. According to [Lub94, Corollary 6.2.1], there exists an invertible power series $a\in\mathscr{S}_{0}(\mathbb{F}_{q})$ satisfying
$f^{\circ r}(X)\bmod(\omega_{L})=a(X^{q}).$
@@
with respect to the identification $X_{n}\mapsto f^{\circ r}(X_{n+1})$ for all integers $n\geqslant 0$, and we denote by $A_{\infty}(\mathcal{O}_{L})$ its $\omega_{L}$-adic completion. We endow this ring with an action of $G_{L}$ by letting $g.X_{n}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=[\chi_{f}(g)]_{f}(X_{n})$ for every integer $n\geqslant 0$.
@@
The ring $A_{\infty}(\mathcal{O}_{L})$ is a strict $\omega_{L}$-ring.
@@
By construction, the ring $A_{\infty}(\mathcal{O}_{L})$ is $\omega_{L}$-adically complete, and $\omega_{L}$ is not a zero divisor. To show that $A_{\infty}(\mathcal{O}_{L})/(\omega_{L})$ is a perfect ring of characteristic $p$, observe that
@@
Endowed with the $(\omega_{L},X_{0})$-adic topology, every formal variable $X_{n}$ is topologically nilpotent in $A_{\infty}(\mathcal{O}_{L})$. Moreover, this ring is universal for $f^{\circ r}$-consistent sequences in the following sense: given a pair $(R,s)$, where $R$ is an $\omega_{L}$-adically complete, ind-complete adic $\mathcal{O}_{L}$-algebra and $s=(s_{n})_{n\geqslant 0}$ is a $f^{\circ r}$-consistent sequence of topologically nilpotent elements in $R$, there exists a unique homomorphism $\phi_{s}\colon A_{\infty}(\mathcal{O}_{L})\rightarrow R$ sending each $X_{n}$ to $s_{n}$ for every $n\geqslant 0$. Conversely, any homomorphism from $A_{\infty}(\mathcal{O}_{L})$ to $R$ arises in this way for some sequence $s$.
@@ 2.4. Regularity of the Galois action
$\vartheta_{f}\colon T_{f}\longrightarrow\mathcal{R}\colon(x_{n})_{n\geqslant 0}\longmapsto(a^{\circ n}(x_{nr}\bmod\omega_{L}))_{n\geqslant 0}.$
Then we apply the universal mapping property of the ring $A_{\infty}(\mathcal{O}_{L})$ to the pair $(\mathcal{R},\varpi)$, where $\varpi=(\varpi_{n})_{n\geqslant 0}$ is the $f^{\circ r}$-consistent sequence of topologically nilpotent elements in $\mathcal{R}$ given by $\varpi_{n}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=(a^{\circ m}(\pi_{(n+m)r}\bmod\omega_{L}))_{m\geqslant 0}$ for each $n\geqslant 0$. Let $\phi_{\varpi}\colon A_{\infty}(\mathcal{O}_{L})\rightarrow\mathcal{R}$ be the corresponding homomorphism. As $\omega_{L}$ is zero in $\mathcal{R}$, this map factors through $(\omega_{L})$; it is even its kernel.
@@
The homomorphism $\phi_{\varpi}$ is $G_{L}$-equivariant with kernel $\omega_{L}A_{\infty}(\mathcal{O}_{L})$.
@@
Proposition 2.4 asserts that $A_{\infty}(\mathcal{O}_{L})$ is a strict $\omega_{L}$-ring, hence the factorization of $\phi_{\varpi}$ through its kernel lifts to an injective $G_{L}$-equivariant homomorphism $A_{\infty}(\mathcal{O}_{L})\hookrightarrow A_{\mathrm{inf},L}$. Under this embedding, we let $x_{n}$ be the image of $X_{n}$ for all integers $n\geqslant 0$. The action of the group $G_{L}$ on these points is particularly interesting, as we shall prove that applying the $f$-logarithm to them produces periods of $\chi_{f}$. As it stands, by [Ber14, Lemma 3.2], we already know that
$t_{f}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{Log}_{f}(x_{0})\in B_{\mathrm{dR}}^{+}.$
This series converges in $B_{\mathrm{dR}}^{+}$ because $\operatorname{Log}_{f}$ converges on $\mathfrak{m}_{\mathbb{C}_{p}}\!$ and $\theta(x_{n})\equiv\pi_{nr}\bmod(\omega_{L})$ for all $n\geqslant 0$. Considering $\operatorname{Log}_{f}(x_{n})$ for a different integer $n>0$ would yield a $\mathbb{Q}_{p}$-multiple of $t_{f}$. Note that $t_{f}$ depends not only on $f$ but also on the choice of the consistent sequence $\pi$ in subsection 2.1. We proceed to show that it is a nonzero period of $\chi_{f}$.
@@
The previous convergence in $B_{\mathrm{dR}}^{+}$ is with respect to the $\ker(\theta)$-adic topology. So there exists some $m\geqslant 1$ for which $\theta(b_{m}x_{0}^{m})=b_{m}\theta(x_{0})^{m}=0$, implying $\theta(x_{0})=0$, where $b_{m}\in K^{\times}$ is the $m$-th coefficient of the formal power series $\operatorname{Log}_{f}$. Consequently, $\theta(t_{f})=0$ and so $\chi_{f}$ has weight $\geqslant 1$. If it had weight $2$, then the $\upsilon_{\mathcal{R}}$-valuation of $x_{0}\bmod(\omega_{L})=\varpi_{0}$ would be at least $2$, however
@@
is less than $2$ when $p$ is odd, a contradiction. For $p=2$, it suffices to compute the next component of $x_{0}$ and this leads to a contradiction too. In particular $t_{f}$ is nonzero. For every $g\in G_{L}$, we have
@@
and hence $\chi_{f}$ is de Rham of Hodge-Tate weight $1$. To show that $\chi_{f}$ is indeed crystalline, it suffices to observe that $t_{f}$ lies in $L\otimes_{L_{0}}B_{\mathrm{cris}}^{+}$ because $\theta(x_{0})=0$ and $\operatorname{Log}_{f}\in\mathscr{S}_{0}(K)$, and its existence implies the existence of a crystalline period since
@@
While in [Spe18] the period $t_{f}$ is already in $B_{\mathrm{cris}}^{+}$, this is not the case here in general. A major obstruction arises from the fact that the formal power series $\operatorname{Log}_{f}$ is not, a priori, defined over the unramified part of $K/\mathbb{Q}_{p}$ when $e>1$.
@@ 2.5. Defining the Tate module structure
Let us summarize the situation up to this point. Starting from $(f,u)$, we have, over a finite extension $L/K$, recovered a restriction of the action of $G_{K}$ on cells of $T_{f}$ via the one of the $\mathbb{Z}_{p}$-iterates of $u$. The chosen resulting character $\chi_{f}$ of $G_{L}$ is such that $\mathbb{Q}_{p}(\chi_{f})$ is crystalline of Hodge-Tate weight $1$, and thus encodes the data of a formal group. Nevertheless, this construction severs the direct link with the initial set $T_{f}$. Our final goal for this section is to relate the information carried by $\mathbb{Z}_{p}(\chi_{f})$ back to the $G_{L}$-set $T_{f}$, thereby endowing it with a $\mathbb{Z}_{p}$-module structure that enables the recovery of the latent formal group.
@@ 2.5.1. A pivotal formal group
Since $\chi_{f}$ is crystalline of Hodge-Tate weight $1$, there exists a height-$1$ formal group $H$ over $\mathcal{O}_{L}$ such that $\operatorname{T}_{p}(H)\simeq\mathbb{Z}_{p}(\chi_{f})$, as stated in §2.2.5. This formal group may not be the one we are looking for, but if Lubin’s conjecture holds true, it would be isomorphic to it. We will transport the structure of $\operatorname{T}_{p}(H)$ onto $T_{f}$. Notice that any other choice of formal group $H^{\prime}$ such that $\operatorname{T}_{p}(H^{\prime})\simeq\mathbb{Z}_{p}(\chi_{f})$ would induce the same structure on $T_{f}$.
@@
The action of $G_{L}\!$ on $\operatorname{T}_{p}(H)$ is given by $\chi_{f}$.
@@
Let $\chi$ be the character of $\operatorname{T}_{p}(H)$. Since there exists a $\mathbb{Z}_{p}[G_{L}]$-module isomorphism $\tau\colon\!\operatorname{T}_{p}(H)\rightarrow\mathbb{Z}_{p}(\chi_{f})$, for every $x\in\operatorname{T}_{p}(H)$ and all $g\in G_{L}$, one has $\chi(g)\tau(x)=\chi_{f}(g)\tau(x)$. This implies that $\chi=\chi_{f}$ given that $\mathbb{Z}_{p}(\chi_{f})$ is free over $\mathbb{Z}_{p}$. ∎
We certainly wish $(f,u)$ to remain a pair of endomorphisms of $T_{f}$ after transporting the structure. To this end we identify the unique endomorphisms of $H$ to which $f$ and $u$ should correspond by examining how they act on the constructed period $t_{f}$, which is a shared data. Indeed, $t_{f}$ was constructed from the dynamical system and by Lemma 2.8 it is also a period of $\operatorname{T}_{p}(H)$. Let $s\in\{f,u\}$; therefore $s$ acts naturally on $T_{f}$. We want to define $s(t_{f})$ arising from this action in a way compatible with the construction of $t_{f}$. Recall that in §2.4 we have followed this path in order to obtain a period:
where the dashed line is defined for all $x\in A_{\mathrm{inf},L}\cap\ker(\theta)$. Considering maps in the same order as we did previously, $s(\pi)$ is sent to the element $(a^{\circ m}(s(\pi_{mr})\bmod\omega_{L}))_{m\geqslant 0}$ which is well defined in $\mathcal{R}$ because $s\bmod(\omega_{L})$ commutes with $a$. This element is nothing but $s(\varpi_{0})$ and is the image of $s(X_{0})$ under $\phi_{\varpi}$. Hence these maps agree after action of $s$. By functoriality of the Witt vectors functor, $s(X_{0})$ is sent to $s(x_{0})$ and lifts $s(\varpi_{0})$ modulo $(\omega_{L})$. As $s(x_{0})\in\ker(\theta)$, we let
$s(t_{f})\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{Log}_{f}(s(x_{0}))=s^{\prime}(0)t_{f}.$
This definition is natural and coincides with the action of a formal group endomorphism on its period, as discussed in paragraph 2.2.5. It thus establishes how $f$ and $u$ act on $t_{f}$. From the perspective of $H$, we assert that $s^{\prime}(0)t_{f}$ arises from $s_{H}$, the multiplication-by-$s^{\prime}(0)$ endomorphism of $H$. Indeed, the composition of the functors considered in §2.2 is faithful:
$\operatorname{\mathbf{Rep}}_{\mathbb{Q}_{p}}^{\mathrm{cris},1}(G_{L})$$\operatorname{\mathbf{MF}}_{L}^{\mathrm{adm},1}(\varphi)$$\operatorname{\mathbf{FGL}\!}/\mathcal{O}_{L}$$\operatorname{\mathbf{Rep}}_{\mathbb{Z}_{p}}^{\mathrm{cris},1}(G_{L})$$\scriptstyle{\operatorname{\mathbf{D}}^{*}_{\mathrm{cris}}}$$\scriptstyle{\operatorname{T}_{p}}$$\scriptstyle{\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}-}$
and $s_{H}$ acts in the same way on all de Rham periods of $\operatorname{V}_{\!p}(H)$, including the crystalline ones. We must therefore associate $f$ with $f_{H}$ when constructing a bijection from $T_{f}$ to $\operatorname{T}_{p}(H)$. We also want this bijection to be $G_{L}$-equivariant. However, Lemma 2.8 implies that it is the $\mathbb{Z}_{p}$-iterates of $u_{H}$ that give the global action of $G_{L}$ on $\operatorname{T}_{p}(H)$. This requires us to take into account the weak knowledge (2) in order to preserve the $G_{L}$-action on $T_{f}$, although it does not guarantee that $u$ is associated to $u_{H}$ and hence that it becomes an endomorphism of $T_{f}$ after transport of structure. Nevertheless, it suffices that $f$ be a formal group endomorphism for $u$ to be one as well.
2.5.2. Transport of structure
Given that $H$ has height $1$, its group of points in $\mathfrak{m}_{\mathbb{C}_{p}}\!$ has exactly $p^{n}$ points of $p^{n}$-torsion, which are precisely the roots of $f^{\circ n}_{H}$, for all integers $n\geqslant 0$. To facilitate the construction of a bijection from $T_{f}$ to $\operatorname{T}_{p}(H)$, we reinterpret the latter Tate module by taking the transition maps to be $f_{H}$ instead of $p]_{H}$, which is legitimate because $f^{\prime}(0)$ is a uniformizer in $\mathbb{Z}_{p}$. Considering the above facts, it is clear that such a bijection exists. We now proceed to construct one satisfying our constraints.
Proposition 2.9.
There exists a $\mathbb{Z}_{p}[G_{L}]$-module structure on the set $T_{f}\!$ for which it is a crystalline character of $G_{L}\!$ of Hodge-Tate weight $1$, and $f\in\operatorname{End}_{\mathbb{Z}_{p}[G_{L}]}(T_{f})$.
Proof.
@@ 2.5.2. Transport of structure
Analogously to the constructions in §2.1, we define $S_{H}\subset\operatorname{T}_{p}(H)$ to be the subset of elements with a nonzero second entry. Given $\beta=(\beta_{n})_{n\geqslant 0}\in S_{H}$, we let

### 22 Apr 2026 v2

@@ 2.3. The universal $f^{\circ r}$-consistent ring
Denote by $r$ the residue degree of $L/\mathbb{Q}_{p}$, so that $q=p^{r}$. To suit this parameter, we adapt the universal $f$-consistent ring in [Spe18] by replacing $f$ with its $r$-th iterate. According to [Lub94, Corollary 6.2.1], there exists an invertible power series $a\in\mathscr{S}_{0}(\mathbb{F}_{q})$ satisfying
$f^{\circ r}(X)\bmod(\pi_{L})=a(X^{q}).$
@@
with the identification $X_{n}\mapsto f^{\circ r}(X_{n+1})$ for all integer $n\geqslant 0$. We let $A_{\infty}(\mathcal{O}_{L})$ be its $\pi_{L}$-adic completion and equip it with a $G_{L}$-action given by $g.X_{n}=[\chi_{f}(g)]_{f}(X_{n})$.
@@
The ring $A_{\infty}(\mathcal{O}_{L})$ is a strict $\pi_{L}$-ring.
@@
The ring $A_{\infty}(\mathcal{O}_{L})$ is a strict $\pi_{L}$-ring if it is $\pi_{L}$-adically complete, $\pi_{L}$-torsion free and $A_{\infty}(\mathcal{O}_{L})/(\pi_{L})$ is a perfect ring. The first two conditions are satisfied by definition. To show the third one, observe that
$A_{\infty}(\mathcal{O}_{L})/(\pi_{L})$ $\simeq A_{\infty}^{\circ}(\mathcal{O}_{L})/(\pi_{L})$
[…]
@@
Endowed with the $(\pi_{L},X_{0})$-adic topology, every formal variable $X_{n}$ is topologically nilpotent in $A_{\infty}(\mathcal{O}_{L})$. Furthermore, this ring is universal for $f^{\circ r}$-consistent sequences in the following sense: given a pair $(R,s)$, where $R$ is a $\pi_{L}$-adically complete, ind-complete adic $\mathcal{O}_{L}$-algebra, $s=(s_{n})_{n\geqslant 0}$ is a $f^{\circ r}$-consistent sequence of topologically nilpotent elements in $R$, there is a unique homomorphism $\phi_{s}\colon A_{\infty}(\mathcal{O}_{L})\rightarrow R$ sending $X_{n}$ to $s_{n}$ for all $n\geqslant 0$. Conversely, any homomorphism from $A_{\infty}(\mathcal{O}_{L})$ to $R$ arises in this way for some sequence $s$.
@@ 2.4. Regularity of the Galois action
$\vartheta_{f}\colon T_{f}\longrightarrow\mathcal{R}\colon(y_{n})_{n\geqslant 0}\longmapsto(a^{\circ n}(y_{nr}\bmod\pi_{L}))_{n\geqslant 0}.$
Then we apply the universal property of the ring $A_{\infty}(\mathcal{O}_{L})$ to the pair $(\mathcal{R},\varpi)$, where $\varpi=(\varpi_{n})_{n\geqslant 0}$ is the $f^{\circ r}$-consistent sequence of topologically nilpotent elements in $\mathcal{R}$ given by
$\varpi_{n}=(a^{\circ m}(\pi_{(n+m)r}\bmod\pi_{L}))_{m\geqslant 0}$
for all $n\geqslant 0$. Let $\phi_{\varpi}\colon A_{\infty}(\mathcal{O}_{L})\rightarrow\mathcal{R}$ be the corresponding homomorphism. As $\pi_{L}$ is zero in $\mathcal{R}$, this map factors through $(\pi_{L})$; it is even its kernel.
@@
The previous homomorphism $\phi_{\varpi}$ is $G_{L}$-equivariant with kernel $\pi_{L}A_{\infty}(\mathcal{O}_{L})$.
@@
Proposition 2.4 asserts that $A_{\infty}(\mathcal{O}_{L})$ is a strict $\pi_{L}$-ring, consequently the factorization of $\phi_{\varpi}$ through its kernel lifts to an injective $G_{L}$-equivariant homomorphism $A_{\infty}(\mathcal{O}_{L})\hookrightarrow A_{\mathrm{inf},L}$. Under this embedding, we let $x_{n}$ be the image of $X_{n}$ for every integer $n\geqslant 0$. The action of the group $G_{L}$ on these points is particularly interesting, as we shall prove that applying the $f$-logarithm to them produces periods of $\chi_{f}$. As it stands, by [Ber14, Lemma 3.2], we already know that
$t_{f}=\operatorname{Log}_{f}(x_{0})\in B_{\mathrm{dR}}^{+}.$
This series converges in $B_{\mathrm{dR}}^{+}$ since $\operatorname{Log}_{f}$ converges on $\mathfrak{m}_{\mathbb{C}_{p}\!}$ and $\theta(x_{n})\equiv\pi_{nr}\bmod(\pi_{L})$ for all $n\geqslant 0$. Considering $\operatorname{Log}_{f}(x_{n})$ for a different integer $n>0$ would yield a $\mathbb{Q}_{p}$-multiple of $t_{f}$. Note that $t_{f}$ depends not only on $f$ but also on the choice of the consistent sequence $\pi$ in Subsection 2.1. We proceed to show that it is a nonzero period of $\chi_{f}$.
@@
The previous convergence in $B_{\mathrm{dR}}^{+}$ is with respect to the $\ker(\theta)$-adic topology. So there exists some $m\geqslant 1$ for which $\theta(b_{m}x_{0}^{m})=b_{m}\theta(x_{0})^{m}=0$, implying $\theta(x_{0})=0$, where $b_{m}\in K^{\times}$ is the $m$-th coefficient of the formal power series $\operatorname{Log}_{f}$. Therefore $\theta(t_{f})=0$ and hence $\chi_{f}$ has weight at least $1$. If it had weight $2$, the $\upsilon_{\mathcal{R}}$-valuation of $x_{0}\bmod(\pi_{L})=\varpi_{0}$ would be at least $2$, however
@@
is less than $2$ when $p$ is odd, a contradiction. For $p=2$, it suffices to compute the next component of $x_{0}$ and this leads to a contradiction. In particular $t_{f}$ is nonzero. For every $g\in G_{L}$, we have
@@
and hence $\chi_{f}$ is de Rham of Hodge-Tate weight $1$. To show that $\chi_{f}$ is crystalline, first observe that $t_{f}$ lies in $L\otimes_{L_{0}}B_{\mathrm{cris}}^{+}$ because $\theta(x_{0})=0$ and $\operatorname{Log}_{f}^{\prime}\in\mathcal{O}_{K}[\![X]\!]$. This implies the existence of a crystalline period since
@@
While in [Spe18] the period $t_{f}$ is in $B_{\mathrm{cris}}^{+}$, this is no more the case here in general. A major obstruction arises from the fact that the formal power series $\operatorname{Log}_{f}$ is not, a priori, defined over the unramified part of $K/\mathbb{Q}_{p}$ when $e>1$.
@@ 2.5. Defining the Tate module structure
Let us summarize the situation up to this point. Starting from $(f,u)$, we have, over a finite extension $L/K$, recovered a restriction of the action of $G_{K}$ on cells of $T_{f}$ via the action of the $\mathbb{Z}_{p}$-iterates of $u$. The chosen resulting $G_{L}$-character $\chi_{f}$ is such that $\mathbb{Q}_{p}(\chi_{f})$ is crystalline of weight $1$, and thus encodes the data of a formal group. Nevertheless, this construction severs the direct link with the initial set $T_{f}$. Our final goal for this section is to relate the information carried by $\mathbb{Z}_{p}(\chi_{f})$ back to the $G_{L}$-set $T_{f}$, thereby endowing it with a $\mathbb{Z}_{p}$-module structure that enables the recovery of the latent formal group.
@@ 2.5.1. A pivotal formal group
Since $\chi_{f}$ is crystalline of Hodge-Tate weight $1$, there exists a height-one formal group $H$ defined over $\mathcal{O}_{L}$ such that
$\operatorname{T}_{p}(H)\simeq\mathbb{Z}_{p}(\chi_{f})$
as stated in Paragraph 2.2.5. This formal group may not be the one we are looking for but, if Lubin’s conjecture holds, it would be isomorphic to it. We will transport the structure of $\operatorname{T}_{p}(H)$ onto $T_{f}$. Any other choice of formal group $H^{\prime}$ over $\mathcal{O}_{L}$ such that $\operatorname{T}_{p}(H^{\prime})\simeq\mathbb{Z}_{p}(\chi_{f})$ would induce the same structure on $T_{f}$. Since $\operatorname{T}_{p}(H)$ and $\mathbb{Z}_{p}(\chi_{f})$ are isomorphic free $\mathbb{Z}_{p}[G_{L}]$-modules of rank $1$, we obtain:
@@
The action of $G_{L}\!$ on $\operatorname{T}_{p}(H)\!$ is given by the character $\chi_{f}\colon G_{L}\rightarrow\mathbb{Z}_{p}^{\times}\!$.
We certainly wish $(f,u)$ to remain a pair of endomorphisms of $T_{f}$ after transporting the structure. To this end, we identify the unique endomorphisms of $H$ to which $f$ and $u$ should correspond by examining how they act on the constructed period $t_{f}$, which is a shared data. Indeed, $t_{f}$ was constructed from the dynamical system and by Lemma 2.8 it is also a period of $\operatorname{T}_{p}(H)$. Let $s\in\{f,u\}$; therefore $s$ acts naturally on $T_{f}$. We want to define $s(t_{f})$ arising from this action in a manner compatible with the construction of $t_{f}$. Recall that in §2.4 we have established this path in order to obtain a period:
where the dashed line is defined for all $x\in A_{\mathrm{inf},L}\cap\ker(\theta)$. Considering maps in the same order as we did previously, $s(\pi)$ is sent to $(a^{\circ m}(s(\pi_{mr})\bmod\pi_{L}))_{m\geqslant 0}$ which is well defined in $\mathcal{R}$ because $s\bmod(\pi_{L})$ commutes with $a$. This element is nothing but $s(\varpi_{0})$ and is the image of $s(X_{0})$ under $\phi_{\varpi}$. Hence these maps agree after action of $s$. By functoriality of the Witt vectors functor, $s(X_{0})$ is sent to $s(x_{0})$ and lifts $s(\varpi_{0})$ modulo $(\pi_{L})$. As $s(x_{0})\in\ker(\theta)$, we let
$s(t_{f})=\operatorname{Log}_{f}(s(x_{0}))=s^{\prime}(0)t_{f}.$
This definition is natural and coincides with the action of a formal group endomorphism on its period, as discussed in Paragraph 2.2.5. It thus establishes how $f$ and $u$ act on $t_{f}$. From the perspective of $H$, we assert that $s^{\prime}(0)t_{f}$ arises from $s_{H}$, the multiplication-by-$s^{\prime}(0)$ endomorphism of $H$. Indeed, the composite of the functors considered in §2.2 is faithful:
$\operatorname{\mathbf{Rep}}_{\mathbb{Q}_{p}}^{\mathrm{cris},\mathrm{fgl}}(G_{L})$$\operatorname{\mathbf{MF}}_{L}^{\mathrm{adm},\mathrm{fgl}}(\varphi)$$\operatorname{\mathbf{FGL}\!}/\mathcal{O}_{L}$$\operatorname{\mathbf{Rep}}_{\mathbb{Z}_{p}}^{\mathrm{cris},\mathrm{fgl}}(G_{L})$$\scriptstyle{\operatorname{\mathbf{D}}^{*}_{\mathrm{cris}}}$$\scriptstyle{\operatorname{T}_{p}}$$\scriptstyle{\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}-}$
where the superscript “fgl” stands for the formal group type condition on the weights or jumps in $\{0,1\}$, and $s_{H}$ acts in the same way on all de Rham periods of $\operatorname{V}_{\!p}(H)$, including the crystalline ones. We must therefore associate $f$ with $f_{H}$ when defining a bijection between $T_{f}$ and $\operatorname{T}_{p}(H)$. We also want this bijection to be $G_{L}$-equivariant. However, Lemma 2.8 implies that the action of $G_{L}$ on $\operatorname{T}_{p}(H)$ is given by the one of the $\mathbb{Z}_{p}$-iterates of $u_{H}$, independently of the cell. Hence we must take into account the weak knowledge (2) in order to be compatible with the existing action of $G_{L}$ on $T_{f}$, although this does not guarantee that $u$ is associated to $u_{H}$ nor that it becomes an endomorphism of $T_{f}$ after transport of structure. Nevertheless, it suffices that $f$ be a formal group endomorphism for $u$ to be one as well.
2.5.2. Transport of structure
Since $H$ has height $1$, its group $H[p^{n}]$ of $p^{n}$-torsion points in $\mathfrak{m}_{\mathbb{C}_{p}\!}$ has exactly $p^{n}$ elements, and these points are also the roots of $f^{\circ n}_{H}$ because $f^{\prime}(0)$ is a uniformizer in $\mathbb{Z}_{p}$. To define a bijection between $T_{f}$ and $\operatorname{T}_{p}(H)$ more easily, we reinterpret the latter Tate module by taking the transition maps to be $f_{H}$ instead of $p]_{H}$:
[…]
which is legitimate since $f^{\prime}(0)$ is a uniformizer in $\mathbb{Z}_{p}$. Considering the above facts, it is clear that such a bijection exists. We now proceed to construct one satisfying our constraints.
Proposition 2.9.
There exists a $\mathbb{Z}_{p}[G_{L}]$-module structure on the set $T_{f}$ for which it is a crystalline $G_{L}$-character of Hodge-Tate weight $1$ and $f\in\operatorname{End}_{\mathbb{Z}_{p}[G_{L}]}(T_{f})$.
@@ 2.5.2. Transport of structure
Analogously to the constructions in §2.1, we define $S_{H}$ to be the subset of $\operatorname{T}_{p}(H)$ of elements with a nonzero second entry. Given $\beta=(\beta_{n})_{n\geqslant 0}\in S_{H}$, let
$S_{H,\beta}=\{(y_{n})_{n\geqslant 0}\in\operatorname{T}_{p}(H)\mid y_{0}=\beta_{0},\dots,y_{\ell}=\beta_{\ell}\}\subset S_{H}$
be a cell. In order to define a $G_{L}$-equivariant bijection $\tau\colon T_{f}\rightarrow\operatorname{T}_{p}(H)$ transporting $f$ to $f_{H}$, we begin by matching $S_{f}$ with $S_{H}$. Set $j=p^{\ell}-p^{\ell-1}$ and choose a system $(\alpha_{1},\dots,\alpha_{j})$ of elements of $S_{f}$ inducing a partition of $S_{f}$ as in §2.1. Similarly, pick a system $(\beta_{1},\dots,\beta_{j})$ of elements of $S_{H}$ where we run through the torsion points of $H$ of exact order $p^{\ell}$. Such a system, in turn, induces a partition of $S_{H}$. For every integers $i\in\{1,\dots,j\}$ and $n\geqslant 0$, and for all $g\in G_{L}$, set
(3) $\tau((f^{\circ n}\circ[\chi_{f,\alpha_{i}}(g)]_{f})(\alpha_{i}))=(f_{H}^{\circ n}\circ[\chi_{f}(g)]_{H})(\beta_{i})$
and $\tau(0)=0$. Fixing $n=0$ and taking $i\in\{1,\dots,j\}$ in (3) furnishes a well-defined bijection $\tau\colon S_{f,\alpha_{i}}\rightarrow S_{H,\beta_{i}}$ since the $\mathbb{Z}_{p}$-iterates of $u$ act simply transitively on $S_{f,\alpha_{i}}$ and the same holds about $u_{H}$ and $S_{f,\beta_{i}}$. Varying $n\geqslant 0$ in (3) allows to reach the remaining points of $T_{f}$ and $\operatorname{T}_{p}(H)$. Therefore (3) defines a $G_{L}$-equivariant bijection $\tau\colon T_{f}\rightarrow\operatorname{T}_{p}(H)$ associating $f$ with $f_{H}$, and the statement follows by transport of structure. ∎
Remark 2.10.
A different choice of representatives for the partitions in the previous proof would provide the same $\mathbb{Z}_{p}[G_{L}]$-module structure on $T_{f}$, but would permute the roles of the individual elements. This is immaterial for our purpose since what matters to us are the group law and the property that the power series $f$ remains an endomorphism of $T_{f}$ after transport of structure, which are unaffected by such choices. Nevertheless, the latent formal group $F$ obtained in Theorem 3.1 will satisfy $\operatorname{T}_{p}(F)=T_{f}$ as $G_{L}$-sets, but only $\operatorname{T}_{p}(F)\simeq T_{f}$ as $\mathbb{Z}_{p}$-modules.
3. The latent formal group
Possessing a $p$-adic Tate module and having access to the existence of a quasi-inverse to the equivalence in Paragraph 2.2.5 is not sufficient to solve Lubin’s conjecture: it is not enough to work up to isomorphism; one must have equality. For this purpose, we make sure that $f$ remains an endomorphism at each step of our constructions, until we obtain a formal group. Recall that $f$ acts on $T_{f}$ by componentwise evaluation or, equivalently, by multiplication by its linear coefficient, and all its other coefficients are uniquely determined by the action of $\mathbb{Z}_{p}$ on $T_{f}$. Our strategy requires an explicit approach in integral $p$-adic Hodge theory. One such approach proceeds through the work of Kisin in [Kis06] and [Kis09] to obtain a formal Breuil-Kisin module, then via the one of Zink in [Zin01] and [Zin02] on displays, indicated by the dashed path in the diagram: